# Quantum Polymorphisms and Commutativity Gadgets

### v1

@@ 2.3. Operations on quantum homomorphisms
We describe here certain constructions taking as input two quantum homomorphisms and giving another quantum homomorphism as the output.
@@
Recall that if $Q$ is a linear operator on $H$ and $Q^{\prime}$ is a linear operator on $H^{\prime}$, then $Q\oplus Q^{\prime}$ is the operator on $H\oplus H^{\prime}$ defined by $(Q\oplus Q^{\prime})(v+v^{\prime})=Q(v)+Q(v^{\prime})$. We extend this definition to quantum functions $Q\colon X\to_{H}A$, $Q^{\prime}:X\to_{H^{\prime}}A$ by letting $Q\oplus Q^{\prime}:X\to_{H\oplus H^{\prime}}A$ be defined by $(Q\oplus Q^{\prime})_{x,a}:=Q_{x,a}\oplus Q^{\prime}_{x,a}$. Note that this is well defined as the set $\{Q_{x,a}\oplus Q^{\prime}_{x,a}\}_{a\in A}$ is a PVM over the Hilbert space $H\oplus H^{\prime}$. Indeed, the fact that its members are projectors onto $H\oplus H^{\prime}$ immediately follows from the definition. Moreover, it is easy to verify that the projector $\sum_{a\in A}(Q_{x,a}\oplus Q^{\prime}_{x,a})$ acts like the identity on both $H$ and $H^{\prime}$—and, hence, on the whole $H\oplus H^{\prime}$ by linearity.
Similarly, if $Q$ is a linear operator on $H$ and $Q^{\prime}$ is a linear operator on $H^{\prime}$, then $Q\otimes Q^{\prime}$ is the linear operator on $H\otimes H^{\prime}$ defined by letting $(Q\otimes Q^{\prime})(v\otimes v^{\prime})=Q(v)\otimes Q^{\prime}(v^{\prime})$ for all $v\in H$ and $v^{\prime}\in H^{\prime}$ and extending by linearity. As for the direct sum, we extend this definition to quantum functions $Q\colon X\to_{H}A$ and $Q^{\prime}\colon X\to_{H^{\prime}}B$ by $(Q\otimes Q^{\prime})_{x,(a,b)}:=Q_{x,a}\otimes Q^{\prime}_{x,b}$ for all $a\in A$ and $b\in B$. This is a quantum function $Q\otimes Q^{\prime}\colon X\to_{H\otimes H^{\prime}}A\times B$.
@@
For $($$\mathbf{QH_{2}}$)), take a symbol $R\in\sigma$ of arity $r$ and let $(x_{1},\dots,x_{r})\in R^{\mathbb{X}}$, $i,j\in\{1,\dots,r\}$, and $a,a^{\prime}\in A$. Then $R_{x_{i},a},R_{x_{j},a^{\prime}}]=[Q_{x_{i},a},Q_{x_{j},a^{\prime}}]\oplus[Q^{\prime}_{x_{i},a},Q^{\prime}_{x_{j},a^{\prime}}]=0$, so $($$\mathbf{QH_{2}}$)) holds. ∎
Observe that classical homomorphisms $h\colon A\to B$ can be identified with quantum homomorphisms $h\colon A\to_{H}B$ over a Hilbert space $H$ of dimension $1$. We will implicitly use this identification in the following. Now, if $Q\colon\mathbb{A}\to_{q}\mathbb{B}$ is such that $Q_{a,b},Q_{a^{\prime},b^{\prime}}]=0$ for all $a,a^{\prime}\in A$ and $b,b^{\prime}\in B$, then $Q=h_{1}\oplus\dots\oplus h_{d}$ for some $d\geq 1$ and $h_{1},\dots,h_{d}\colon\mathbb{A}\to\mathbb{B}$ classical homomorphisms (where $d$ can be taken to be the dimension of $Q$). This is a simple consequence of the spectral theorem.
@@
(1) $Q_{a,b}(x)=\sum_{i=1}^{d}\mathbf{1}[h_{i}(a)=b]\langle x,e_{i}\rangle e_{i}$
for all $x\in H$, where $\mathbf{1}{[h_{i}(a)=b]}$ is $1$ if $h_{i}(a)=b$ and 0 otherwise.
@@
$=R_{x,(a,b)}(v\otimes v^{\prime})$
so that each $R_{x,(a,b)}$ is a projector. Moreover, if $x\in X$, then
$\sum_{a,b}R_{x,(a,b)}$ $=\sum_{a,b}Q_{x,a}\otimes Q_{x^{\prime},b}$
$=\sum_{a}Q_{x,a}\otimes(\sum_{b}Q_{x^{\prime},b})$
$=\sum_{a}Q_{x,a}\otimes\operatorname{id}_{H^{\prime}}$
$=(\sum_{a}Q_{x,a})\otimes\operatorname{id}_{H^{\prime}}=\operatorname{id}_{H}\otimes\operatorname{id}_{H^{\prime}}=\operatorname{id}_{H\otimes H^{\prime}}$
@@
by $(R\bullet Q)_{a,c}=\sum_{b\in B}R_{b,c}\otimes Q_{a,b}$.
Note that the sum in the definition is finite (even if $B$ is infinite), since $H$ has finite dimension and therefore only finitely many of the $Q_{a,b}$ are non-zero.
The fact that the composition of quantum homomorphisms is a quantum homomorphism was proved in [abramsky2017quantum] using that this operation can be expressed as the composition of Kleisli arrows in the Kleisli category for the quantum monad. For the convenience of the reader, we give here a self-contained proof following [abramsky2017quantum].
Lemma 6 ([abramsky2017quantum]).
@@
$S_{a,c}^{\top}=(\sum_{b}R_{b,c}\otimes Q_{a,b})^{\top}=\sum_{b}R_{b,c}^{\top}\otimes Q_{a,b}^{\top}=\sum_{b}R_{b,c}\otimes Q_{a,b}=S_{a,c},$
@@
Note that for every $(b_{1},\dots,b_{r})$, either $(b_{1},\dots,b_{r})\in R^{\mathbb{B}}$—in which case $R_{b_{1},c_{1}}\dots R_{b_{r},c_{r}}=0$ since $R$ satisfies $($$\mathbf{QH_{1}}$))—or $(b_{1},\dots,b_{r})\not\in R^{\mathbb{B}}$—and then $Q_{a_{1},b_{1}}\dots Q_{a_{r},b_{r}}=0$ since $Q$ satisfies $($$\mathbf{QH_{1}}$)). Thus, the sum above evaluates to 0 and $($$\mathbf{QH_{1}}$)) holds.
@@
Let $\mathbb{X}$ be a core and let $Q\colon\mathbb{X}\to_{q}\mathbb{X}$ be a quantum homomorphism. If $Q$ is non-contextual, then
@@
$\sum_{x\in X}Q_{x,y}=\bigoplus_{i\in[d]}\sum_{x\in X}(h_{i})_{x,y}=\operatorname{id}.$
3. Quantum polymorphisms and reductions
In this section, we show that reductions between entangled CSPs can be obtained via a quantised notion of pp-definitions that we call q-definitions. In order to prove that q-definitions exist, we make use of the notion of commutativity gadgets as defined in [Ji]. Then, we prove that the existence of commutativity gadgets is entirely determined by a closure property of the quantum polymorphisms of the given language. We begin by defining q-definitions.

### v2

@@ 3.3 Operations on quantum homomorphisms
In this subsection, we describe certain constructions taking as input two quantum homomorphisms and returning another quantum homomorphism as the output.
@@
Given a linear operator $Q$ on $H$ and a linear operator $Q^{\prime}$ on $H^{\prime}$, we let $Q\oplus Q^{\prime}$ be the linear operator on $H\oplus H^{\prime}$ defined by $(Q\oplus Q^{\prime})(v+v^{\prime})=Q(v)+Q(v^{\prime})$. We extend this definition to quantum functions $Q\colon X\to_{H}A$, $Q^{\prime}\colon X\to_{H^{\prime}}A$ by letting $Q\oplus Q^{\prime}\colon X\to_{H\oplus H^{\prime}}A$ be defined by $(Q\oplus Q^{\prime})_{x,a}:=Q_{x,a}\oplus Q^{\prime}_{x,a}$. Note that this is well defined as the set $\{Q_{x,a}\oplus Q^{\prime}_{x,a}\}_{a\in A}$ is a PVM over the Hilbert space $H\oplus H^{\prime}$. Indeed, the fact that its members are projectors onto $H\oplus H^{\prime}$ immediately follows from the definition. Moreover, it is easy to verify that the projector $\sum_{a\in A}(Q_{x,a}\oplus Q^{\prime}_{x,a})$ acts like the identity on both $H$ and $H^{\prime}$—and, hence, on the whole $H\oplus H^{\prime}$ by linearity. We now show that this operation preserves quantum homomorphisms.
@@
For $($$\mathbf{QH_{2}}$)), take a symbol $R\in\sigma$ of arity $r$ and let $(x_{1},\dots,x_{r})\in R^{\mathbb{X}}$, $i,j\in[r]$, and $a,a^{\prime}\in A$. Then $R_{x_{i},a},R_{x_{j},a^{\prime}}]=[Q_{x_{i},a},Q_{x_{j},a^{\prime}}]\oplus[Q^{\prime}_{x_{i},a},Q^{\prime}_{x_{j},a^{\prime}}]=0$, so $($$\mathbf{QH_{2}}$)) holds. ∎
Observe that classical homomorphisms $h\colon A\to B$ can be identified with quantum homomorphisms $h\colon A\to_{H}B$ over a Hilbert space $H$ of dimension $1$. We will implicitly use this identification in the following. Now, if $Q\colon\mathbb{A}\to_{\operatorname{q}}\mathbb{B}$ is such that $Q_{a,b},Q_{a^{\prime},b^{\prime}}]=0$ for all $a,a^{\prime}\in A$ and $b,b^{\prime}\in B$, then $Q=h_{1}\oplus\dots\oplus h_{d}$ for some $d\geq 1$ and $h_{1},\dots,h_{d}\colon\mathbb{A}\to\mathbb{B}$ classical homomorphisms (where $d$ can be taken to be the dimension of $Q$). This is a simple consequence of the spectral theorem.
@@
$Q_{a,b}(x)=\sum_{i=1}^{d}\mathbf{1}_{[h_{i}(a)=b]}\langle x,e_{i}\rangle e_{i}$ (1)
for all $x\in H$, where $\mathbf{1}_{[h_{i}(a)=b]}$ is $1$ if $h_{i}(a)=b$ and 0 otherwise.
Suppose now that $Q$ is a linear operator on $H$ and $Q^{\prime}$ is a linear operator on $H^{\prime}$, and let $Q\otimes Q^{\prime}$ be the linear operator on $H\otimes H^{\prime}$ defined by setting $(Q\otimes Q^{\prime})(v\otimes v^{\prime})=Q(v)\otimes Q^{\prime}(v^{\prime})$ for all $v\in H$ and $v^{\prime}\in H^{\prime}$ and extending by linearity. As for the direct sum, we extend this definition to quantum functions $Q\colon X\to_{H}A$ and $Q^{\prime}\colon X\to_{H^{\prime}}B$ by $(Q\otimes Q^{\prime})_{x,(a,b)}:=Q_{x,a}\otimes Q^{\prime}_{x,b}$ for all $a\in A$ and $b\in B$. This is a quantum function $Q\otimes Q^{\prime}\colon X\to_{H\otimes H^{\prime}}A\times B$. Like for the direct sum, we now show that the tensor product preserves quantum homomorphisms.
@@
$=R_{x,(a,b)}(v\otimes v^{\prime}),$
so each $R_{x,(a,b)}$ is a projector onto $H\otimes H^{\prime}$. Moreover, if $x\in X$, then
$\sum_{a,b}R_{x,(a,b)}$ $=\sum_{a,b}Q_{x,a}\otimes Q_{x^{\prime},b}=\sum_{a}Q_{x,a}\otimes\sum_{b}Q_{x^{\prime},b}=\operatorname{id}_{H}\otimes\operatorname{id}_{H^{\prime}}=\operatorname{id}_{H\otimes H^{\prime}},$
@@
by $(R\bullet Q)_{a,c}=\sum_{b\in B}R_{b,c}\otimes Q_{a,b}$ for each $a\in A$, $c\in C$.
The fact that the composition of quantum homomorphisms is a quantum homomorphism was proved in [1], and it corresponds to the fact that this operation can be expressed as the composition of Kleisli arrows in the Kleisli category for the quantum monad. (See also [48] for a recent extension of the Kleisli framework to different quantum relaxations of classical morphisms.) For the convenience of the reader, we give here a self-contained proof following [1].
Lemma 21 ([1]).
@@
$S_{a,c}^{*}=(\sum_{b}R_{b,c}\otimes Q_{a,b})^{*}=\sum_{b}R_{b,c}^{*}\otimes Q_{a,b}^{*}=\sum_{b}R_{b,c}\otimes Q_{a,b}=S_{a,c},$
@@
Note that for every $(b_{1},\dots,b_{r})\in B^{r}$, either $(b_{1},\dots,b_{r})\in R^{\mathbb{B}}$—in which case $R_{b_{1},c_{1}}\dots R_{b_{r},c_{r}}=0$ since $R$ satisfies $($$\mathbf{QH_{1}}$))—or $(b_{1},\dots,b_{r})\not\in R^{\mathbb{B}}$—and then $Q_{a_{1},b_{1}}\dots Q_{a_{r},b_{r}}=0$ since $Q$ satisfies $($$\mathbf{QH_{1}}$)). Thus, the sum above evaluates to 0 and $($$\mathbf{QH_{1}}$)) holds.
@@
We also generalise the composition of quantum functions to functions with arity greater than 1.
Definition 22.
Fix positive integers $k_{1},\dots,k_{n}$, and let $W_{1},\dots,W_{n}$ be quantum functions, where $W_{i}\colon A^{k_{i}}\to_{\operatorname{q}}B$ for all $i\in[n]$. Take also a quantum function $Q\colon B^{n}\to_{\operatorname{q}}C$. We define
$Q*(W_{1},\dots,W_{n})\colon A^{\sum k_{i}}\to_{\operatorname{q}}C$
by $(Q*(W_{1},\dots,W_{n}))_{(\mathbf{a}^{(1)},\dots,\mathbf{a}^{(n)}),c}=\sum_{\mathbf{b}\in B^{n}}Q_{\mathbf{b},c}\otimes\bigotimes_{i=1}^{n}(W_{i})_{\mathbf{a}^{(i)},b_{i}}$.
Note that this is a well-defined quantum function, since
$\sum_{c\in C}(Q*(W_{1},\dots,W_{n}))_{(\mathbf{a}^{(1)},\dots,\mathbf{a}^{(n)}),c}$ $=\sum_{c\in C}\sum_{\mathbf{b}\in B^{n}}Q_{\mathbf{b},c}\otimes\bigotimes_{i=1}^{n}(W_{i})_{\mathbf{a}^{(i)},b_{i}}$
$=\operatorname{id}\otimes\bigotimes_{i=1}^{n}\sum_{b\in B}(W_{i})_{\mathbf{a}^{(i)},b}=\operatorname{id}.$
Lemma 23.
If $W_{1},\dots,W_{n}$ are homomorphisms $\mathbb{A}^{k_{i}}\to_{H_{i}}\mathbb{B}$ and $Q\colon\mathbb{B}^{n}\to_{H}\mathbb{C}$, then $Q*(W_{1},\dots,W_{n})$ is a homomorphism $\mathbb{A}^{\sum k_{i}}\to_{H\otimes\bigotimes H_{i}}\mathbb{C}$.
Proof.
Consider the quantum function $W:A^{\sum k_{i}}\to_{\operatorname{q}}B^{n}$ defined as follows. For $\mathbf{a}^{(1)}\in A^{k_{1}},\dots,\mathbf{a}^{(n)}\in A^{k_{n}}$ and $\mathbf{b}\in B^{n}$, we set
$W_{((\mathbf{a}^{(1)},\dots,\mathbf{a}^{(n)}),\mathbf{b})}=\bigotimes_{i\in[n]}(W_{i})_{\mathbf{a}^{(i)},b_{i}}.$
Reasoning as in the proof of Lemma 19, we readily check that $W$ is a quantum homomorphism $\mathbb{A}^{\sum k_{i}}\to_{\bigotimes H_{i}}\mathbb{B}^{n}$, and that
$Q*(W_{1},\dots,W_{n})=Q\bullet W.$
Then, the result follows from Lemma 21. ∎
@@
Let $\mathbb{X}$ be a core and let $Q\colon\mathbb{X}\to_{\operatorname{q}}\mathbb{X}$ be a quantum homomorphism. If $Q$ is non-contextual, then
@@
$\sum_{x\in X}Q_{x,y}=\bigoplus_{i\in[d]}\sum_{x\in X}(h_{i})_{x,y}=\operatorname{id},$
as required. ∎
4 Quantum algebraic reductions
In this section, we start our analysis of reductions between quantum CSPs by lifting Theorem 2 to the quantum setting via the notion of quantum polymorphisms. First of all, in the next subsection, we collect information on the algebraic structure of the quantum polymorphism set of a constraint language.
4.1 The algebraic structure of quantum polymorphisms
The set of classical polymorphisms of a relational structure $\mathbb{A}$ can be endowed with the algebraic structure of a function minion over $A$. A function minion over some set $S$ is a subset $\mathscr{M}$} of the set of functions $\{f:S^{n}\to S:n\in\mathbb{N}\}$ that is closed under identifying variables, permuting variables, and adding dummy variables. This can be expressed compactly as follows: For each $n,m\in\mathbb{N}$, each map $\pi:[n]\to[m]$, and each $f\in{\mathscr{M}}$ of arity $n$ (meaning that $f$ is a function $S^{n}\to S$), the function $f_{/\pi}:S^{m}\to S$ defined by
$f_{/\pi}(x_{1},\dots,x_{m})=f(x_{\pi(1)},x_{\pi(2)},\dots,x_{\pi(n)})$
belongs to $\mathscr{M}$}. It is clear from the definition of classical polymorphisms that the set $\operatorname{Pol}(\mathbb{A})$ is invariant under this type of operations (known as minors), so $\operatorname{Pol}(\mathbb{A})$ is indeed a function minion over $A$. In fact, one can endow $\operatorname{Pol}(\mathbb{A})$ with the richer structure of a (function) clone, due to the fact that polymorphisms are closed under composition (and composition is associative). Nevertheless, it was observed in [10] that the complexity of CSPs is entirely determined by the height-one or minor identities holding in $\operatorname{Pol}(\mathbb{A})$; i.e., identities of the form $f=g_{/\pi}$. In particular, in order to capture reducibility of CSPs via pp-definitions and pp-constructions, it is sufficient to compare their polymorphism clones in terms of minor-preserving maps; i.e., so-called minion homomorphisms. The minion framework proved robust enough to capture reducibility and complexity of variants of the CSP; in particular, the promise CSP [3, 19, 5], whose polymorphisms do not compose and, thus, do not form clones. In addition, minions were shown to be well-suited to capture the power of relaxations of (promise) CSPs via multiple algorithmic models [5, 27, 28, 29, 25, 17, 30, 11, 55]; see also [16] for a recent application to CSP sparsification and non-redundancy. In these works, the notion of function minions was gradually lifted to the categorical concept of abstract minions, retaining the algebraic structure of function minions without the restriction of being applied to sets of functions; see, in particular, [18, 6].